OpenStudy (anonymous):

\[\int _a^b \frac{dx}{y}\],where \[ y^2=ax^2+bx+c\]

OpenStudy (experimentx):

reduce y in to this from (px + q)^2 + r

OpenStudy (anonymous):

where \[y^2=ax^2+bx+c\]

OpenStudy (anonymous):

sry i meant y^2

OpenStudy (experimentx):

do the same .. it wouldn't make any difference

OpenStudy (anonymous):

\[1/a(x+b/2a)^2-b^2/4a+c/a

OpenStudy (anonymous):

\[1/a(x+b/2a)^2-b^2/4a+c/a\]

OpenStudy (anonymous):

i tried a euclidian sub of t \[t=y+x\sqrt a\]

OpenStudy (experimentx):

yeah .. then substitute u = Sqrt(a)(x + b/x) should be of the form